Morphine is a painkiller with a chemical formula C17H19NO3 . Suppose 30 mg of morphine is diluted in 100 mL of water .Water has a density of 1g/ mL .  Calculate the average molecular weight of the solution.
Answer #1

First we calculate the moles of each compound in solution:

Morphine: 30 mg * (1 g / 1000 mg) * (1 mol / 285.34 g) = 1.05x10 ^ -4 mol

Water: 100 mL * (1 g / 1 mL) * (1 mol / 18.02 g) = 5.549390 mol

The average molecular weight is calculated with the sum of the moles by the molar weight of each compound, on the sum of the moles in solution:

MM ave = (0.000105 mol * 285.34 g / mol) + (5.549390 mol * 18.02 g / mol) / 5.549390 + 0.000105 mol = 18.025058 g / mol

If you liked this answer, please rate it in a positive way, you would help me a lot, thank you.